OpenStudy (anonymous):

I m having a hard time with this question. I cant figure out what the function would be for Plan A or Plan B You get hired for a new job that will only last for one month (31 days, you work 7 days a week) and you are offered the two following payment plans to choose from: • Plan A: $4,000 per day for the whole month with a $10,000 signing bonus. • Plan B: $0.01 for day 1 $0.02 for day 2 $0.04 for day 3 $0.08 for day 4 $0.16 for day 5 ….and so on every day is double the previous days salary, with no signing bonus. Create two functions that model each of these payment plans, wher

OpenStudy (anonymous):

Okay, let's work on Plan A

OpenStudy (anonymous):

notice that 4000 happen every day whereas 10,000 happen only once

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so let's denote x for a day

OpenStudy (anonymous):

Plan A 4000x+10,000

OpenStudy (anonymous):

Do you understand how I got that before we move on to Plan B

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

Okay, let's move on to Plan B

OpenStudy (anonymous):

The first day you make .01

OpenStudy (anonymous):

second day you make .01*(2)=.02 third day you make .01*(2)*(2)=.04

OpenStudy (anonymous):

So notice for each day we are multiplying by 2

OpenStudy (anonymous):

yes it doubles each day

OpenStudy (anonymous):

second day you make .01*(2)=.02 third day you make .01*(2)*(2)= .01*(2)^2=.04 fourth day you make .01(2)*2*2=.01*(2)^3=.08

OpenStudy (anonymous):

Does that make sense

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes that does

OpenStudy (anonymous):

Okay , so you need to add what you make each day for whole month(31 days)

OpenStudy (anonymous):

.01(2)^0+.01(2)^1+.01(2)^2+........+.01(2)^30 First day second third 31st day

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You can add up all the term by hand, but it is very tedious. We can instead use a formula called Sum of Geometric Series Our terms- .01(2^0+2^1+......+2^30) r-2 n-30 Geometric Series Sum= \[[1-r^{n+1}]\over(1-r)\]

OpenStudy (anonymous):

that's fraction

OpenStudy (anonymous):

ok, it is a long process by hand

OpenStudy (anonymous):

that's why we use geometric series

OpenStudy (anonymous):

ok which numbers do I enter in

OpenStudy (anonymous):

.01 * (1-(2)^31)/(1-2)

OpenStudy (anonymous):

that seems to be quit complicated

OpenStudy (anonymous):

What is?

OpenStudy (anonymous):

I think I figured out how much the salary would be be in 31 days, $21474836

OpenStudy (anonymous):

you are missing two digit at the end 2147483647

OpenStudy (anonymous):

$21474836.48

OpenStudy (anonymous):

* sorry, you are right, can't see the period

OpenStudy (anonymous):

ok what is the function for plan B and here is more information for this question (Plan A’s function, , will represent what you earn total over the whole 31 days, Plan B’s function, , will represent what you earn each individual day.)

OpenStudy (anonymous):

Plan B function for each individual day =nth day= .01*(2)^(n-1)

OpenStudy (anonymous):

Plan A's function A(t) and PLan B's function B(t),

OpenStudy (anonymous):

ok so my question is how would i figure out the salary using that equation for plan B

OpenStudy (anonymous):

add up each day